how much more or less will a passenger pay if she travels five times in April if demand 6400 and price is 18 and 3 times in july if demand is 2000 and price 40?


1
Expert's answer
2021-09-07T12:07:47-0400

"D1\u00d7P1\u00d7T1=D2\u00d7P2\u00d7T2"

"T2\u00d76400\u00d718=2000\u00d740\u00d73"


"115200 T2 = 240000"


"T2 = \frac{240000}{115200}"


"T2 = 2.08"


When rounded off 2.08 to a whole number = 2


"= 3-2 = 1"